A gas with a constant pressure of 150 kPa expands from a volume of 0.76 m3 to a volume of 0.92 m3. How much work does the gas do?


1
Expert's answer
2022-03-09T10:33:25-0500

The work is given as follows:


"W = p\Delta V - p(V_2-V_1)"

where "p = 150\times 10^3Pa", "V_2 = 0.92m^3, V_1 = 0.76m^3". Thus, obtain:


"W = 150\times 10^3Pa\cdot (0.92m^3-0.76m^3) = 24\times 10^3J = 24kJ"

Answer. 24kJ.